### v3

@@ Abstract
0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: Benchmarking LLMs for Extracting and Structuring Financial Information
Yan Wang1, Lingfei Qian1, Xueqing Peng1, Yang Ren1, Yi Han2, Keyi Wang3, Dongji Feng4, Fengran Mo5, Shengyuan Lin6, Qinchuan Zhang7, Kaiwen He3, Chenri Luo3, Jianxing Chen3, Junwei Wu3, Chen Xu1, Ziyang Xu1, Jimin Huang1thanks:  Corresponding author, Guojun Xiong8, Xiao-Yang Liu3, Qianqian Xie1, Jian-Yun Nie5 1The Fin AI, 2Georgia Institute of Technology, 3Columbia University, 4California State University 5University of Montreal, 6Carnegie Mellon University, 7Rensselaer Polytechnic Institute, 8Harvard University Correspondence: jimin.huang@thefin.ai
1 Introduction
@@
Although LLMs excel at question answering qian2025fino1, information extraction balasubramanian2025leveraging, and long-document summarization yuan2025strucsum, their ability to perform financial tagging remains underexplored. Unlike open-ended generation, tagging requires models to commit to a precise semantic interpretation for each reported number, grounding it in dispersed evidence across tables and surrounding text, and aligning it to a standardized concept that downstream systems can interpret consistently. Each year, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, providing essential information for investors, regulators, and analysts. However, inconsistent terminology complicates reliable interpretation. XBRL, introduced in 1999, defines the global standard for machine-readable financial reporting richards2006introduction and is now adopted in 65 countries [^3] https://www.xbrl.org/the-standard/what/what-is-xbrl/. In practice, XBRL tagging requires grounding thousands of reported numbers per filing in mixed table and text evidence and aligning them to over 17,000 standardized taxonomy concepts(Figure 1), a workflow that remains largely manual and error-prone [^4] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
@@
Figure 1: An example of Financial Tagging in realistic scenarios, where the blue numbers mark facts to be tagged.
@@
Nevertheless, current XBRL tagging benchmarks often emphasize what is easiest to annotate rather than what is hardest to understand. Many reduce tagging to classification over a truncated taxonomy and remove table structure that carries substantial semantic evidence in real filings (Table 1). As a result, existing datasets typically cover only a few thousand common concepts instead of the full 17,000+ taxonomy, which limits evaluation on rare but important concepts and makes it difficult to assess generalization under long-tailed label distributions. In addition, prior work commonly formulates tagging as closed-set extreme classification loukas2022finer; sharma2023financial. This framing places heavy weight on recalling label identifiers and becomes brittle when the full taxonomy cannot be represented in-context, increasing the likelihood of invalid or hallucinated tags. Consistent with this concern, our ablations show near-zero performance even for strong LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt under full-taxonomy settings. Finally, existing benchmarks focus primarily on narrative text and omit tables, despite tables being the main carrier of numerical disclosures. Consequently, they do not evaluate the joint reasoning over multi-level tabular structure and surrounding text that has been studied in financial QA chen2021finqa; zhu2021tat and is required for realistic end-to-end tagging.
@@
Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “Num.Reasoning” indicates the numerical reasoning. “Struct.IE” denotes the structured information extraction.
@@
FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,688
FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,688
@@
To address these issues, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-oriented benchmark that evaluates whether models can recover the meaning of financial facts by extracting numerical values from mixed table and text inputs and aligning them to the full US-GAAP taxonomy. Unlike prior work loukas2022finer; sharma2023financial that formulates tagging as flat classification over a truncated label set, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires end-to-end reasoning over both narrative disclosures and tables, and scales evaluation to the complete 17k+ concept space. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a two-stage task formulation that separates decisions typically conflated in extreme classification. Financial Numerical Identification (FinNI) asks models to identify numerical facts together with their supporting evidence, while Financial Concept Linking (FinCL) maps each fact to the correct US-GAAP concept by semantic comparison with taxonomy definitions, reframing tagging as alignment rather than label recall. We release two expert-verified evaluation sets, FinNI-eval and FinCL-eval, derived from real-world XBRL filings, along with a unified protocol that evaluates both extraction accuracy and concept-alignment performance.
Across 13 state-of-the-art LLMs, our results show a consistent extraction–linking gap. While top models reliably recover numerical facts from mixed table and text inputs, they struggle to assign the correct US-GAAP concepts, indicating that representations sufficient for extraction do not consistently encode the latent semantic conditions that distinguish closely related concepts, such as scope, aggregation level, and qualifying context. As a result, FinNI performance is strong, whereas FinCL remains below 0.19 even after candidate retrieval, with most errors arising from near-neighbor concepts that share similar surface cues but differ in these underlying conditions. Pipeline controls further clarify this failure mode: accuracy improves when gold entities are provided and improves again when the gold concept is included among candidates, yet it does not converge, showing that errors persist even when evidence is localized and the correct concept is available. The restricted cash example illustrates this pattern, where the model selects a narrower concept triggered by a salient local phrase, despite broader contextual evidence supporting an aggregate concept.
@@
We conclude our main contributions as follows: (1) A New Formulation: We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that reframes XBRL tagging into a two-stage “Extract-and-Align” paradigm, enabling full-scope evaluation against 17k+ US-GAAP concepts, a scale previously intractable for zero-shot LLMs. (2) High-Fidelity Resources: We release two expert-verified datasets, FinNI-eval and FinCL-eval, which incorporate complex tabular structures to assess joint reasoning across structured and unstructured financial modalities. (3) Diagnostic Insights: By benchmarking 13 SOTA LLMs, we identify a critical “knowledge-alignment gap”, demonstrating that while LLMs excel at numeric extraction, they require our alignment framework to prevent performance collapse during large-scale taxonomy mapping.
2 Related Work
@@
From XBRL Classification to Alignment. Traditional XBRL tagging has been framed as extreme multi-class classification. Benchmarks such as FiNER loukas2022finer and FNXL sharma2023financial evaluate models on truncated label sets (139–2,800 tags), and methods like GalaXC saini2021galaxc, SECBERT loukas2022finer, and AttentionXML sharma2023financial focus on coping with label sparsity through hierarchical modeling. However, these closed-set formulations overlook the semantic reasoning required to navigate the full 17k+ US-GAAP taxonomy and are poorly matched to LLM capabilities. Recent LLM-based tools such as XBRL-Agent han2024xbrlagent explore filing analysis but do not evaluate end-to-end fact-to-taxonomy alignment. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: addresses this gap by reframing tagging as a two-stage extract-and-align task.
Financial Document Understanding. In parallel, financial NLP has progressed from entity and relation extraction (e.g., FiNER-ORD shah2023finer, FinRED sharma2022finred) to numerical reasoning over documents. Benchmarks such as FinQA chen2021finqa, TAT-QA zhu2021tat, and DocVQA mathew2021docvqa require joint reasoning over text and tables. Yet these tasks are designed for QA and do not constrain outputs to standardized accounting concepts. Our benchmark instead requires both multi-modal fact extraction and precise alignment to official US-GAAP definitions, focusing on the “last mile” of structured financial reporting.
3 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
3.1 Task Formulation
@@
Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g. “us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType). Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping. Specifically, monetaryItemType refers to a financial amount, such as revenue or expenses, typically reported in currency units (e.g., USD). percentItemType represents a rate or ratio, expressed as a decimal (e.g., 25%). sharesItemType indicates the number of shares, such as stocks held or issued. perShareItemType captures values reported on a per-share basis, like earnings per share. Lastly, integerItemType is used for whole number counts, such as the number of employees or transactions.
We formalize the task as a mapping:
$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
@@
Inspired by information extraction and alignment works wu2019scalable; wadden2019entity; mukherjee2022ectsum; wang2022conditional and after discussing with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
Financial Numeric Identification (FinNI)
The first subtask of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: focuses on identifying numerical values in a financial document and assigning each a coarse-grained value data type. This corresponds to detecting the Fact and Type components of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task.
We formalize this subtask as a mapping:
$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)

### v4

@@ Abstract.
XBRL tagging, Benchmark, Large language model, Information extraction, Information retrieval, Reranking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Information extractionccs: Information systems Information extractionccs: Information systems Top-k retrieval in databases
1. Introduction
Although LLMs excel at question answering (Qian et al., 2025), information extraction (Balasubramanian et al., 2025), and long-document summarization (Yuan et al., 2025), their ability to perform financial tagging remains underexplored. Unlike open-ended generation, tagging requires committing each reported number to a precise semantic interpretation, grounding it in mixed table and text evidence, and aligning it to standardized concepts interpretable by downstream systems.
@@ 1. Introduction
Each year, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities. However, inconsistent terminology complicates reliable interpretation. XBRL, introduced in 1999, defines the global standard for machine-readable financial reporting (Richards et al., 2006) and is now adopted in 65 countries [^3] https://www.xbrl.org/the-standard/what/what-is-xbrl/ . In practice, XBRL tagging requires grounding thousands of reported numbers per filing in mixed evidence and aligning them to over 17,000 standardized taxonomy concepts (Figure 1), a process that remains largely manual and error-prone [^4] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
@@
Figure 1. An example of Financial Tagging in realistic scenarios, where the blue numbers mark facts to be tagged.
@@
Nevertheless, existing XBRL tagging benchmarks emphasize annotation convenience over semantic difficulty. Many reduce tagging to classification over truncated taxonomies and remove tabular structure that carries critical evidence in real filings (Table 1). As a result, evaluation is limited to a small subset of common concepts, obscuring long-tailed generalization. Moreover, prior work typically formulates tagging as closed-set extreme classification (Loukas et al., 2022; Sharma et al., 2023), which becomes brittle when the full taxonomy cannot be represented in-context and increases the risk of invalid or hallucinated tags. Our ablations confirm this limitation: even strong LLMs show near-zero performance under full-taxonomy settings. In addition, most benchmarks omit tables entirely, despite tables being the primary carrier of numerical disclosures, and thus fail to evaluate the joint reasoning over structured tables and narrative context required in realistic tagging (Chen et al., 2021; Zhu et al., 2021).
@@
Table 1. Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “Num.Reasoning” indicates the numerical reasoning. “Struct.IE” denotes the structured information extraction.
@@
FinQA (Chen et al., 2021) Decision making SEC 10-K QA text/table 0 0
ConvFinQA (Chen et al., 2022) Decision making SEC 10-K QA text/table 0 0
TAT-QA (Zhu et al., 2021) Financial analysis Chinese financial reports QA text/table 0 0
DocVQA (Mathew et al., 2021) Enterprise automation Financial document images QA text/image 0 0
FiNER-ORD (Shah et al., 2023) Financial tagging Chinese financial disclosures Classification text 49 0
FinRED (Sharma et al., 2022) Knowledge graph construction English financial news Classification text 38 0
FiNER (Loukas et al., 2022) XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,688
FNXL (Sharma et al., 2023) XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,688
@@
To address these issues, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-oriented benchmark that evaluates end-to-end financial tagging over mixed table and text inputs with alignment to the full US-GAAP taxonomy. Unlike prior work (Loukas et al., 2022; Sharma et al., 2023), 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: reframes tagging as an extract-and-align process rather than flat classification, enabling evaluation across all 17k+ concepts. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a two-stage formulation. Financial Numerical Identification (FinNI) focuses on extracting numerical facts with supporting evidence, while Financial Concept Linking (FinCL) maps each fact to the correct US-GAAP concept via semantic alignment with taxonomy definitions. We release two expert-verified evaluation sets, FinNI-eval and FinCL-eval, derived from real-world XBRL.
@@
We conclude our main contributions as follows: (1) A New Formulation: We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that reframes XBRL tagging into a two-stage “Extract-and-Align” paradigm, enabling full-scope evaluation against 17k+ US-GAAP concepts, a scale previously intractable for zero-shot LLMs. (2) High-Fidelity Resources: We release two expert-verified datasets, FinNI-eval and FinCL-eval, which incorporate complex tabular structures to assess joint reasoning across structured and unstructured financial modalities. (3) Diagnostic Insights: By benchmarking 13 state-of-the-art LLMs, we identify a critical “knowledge-alignment gap”, demonstrating that while LLMs excel at numeric extraction, they require our alignment framework to prevent performance collapse during large-scale taxonomy mapping.
2. Related Work
@@
Prior XBRL tagging research primarily formulates the task as extreme multi-class classification over truncated label sets, as in FiNER and FNXL (Loukas et al., 2022; Sharma et al., 2023), with methods such as GalaXC, SECBERT, and AttentionXML (Saini et al., 2021; Loukas et al., 2022; Sharma et al., 2023) addressing label sparsity via hierarchical modeling. However, these closed-set formulations do not reflect the semantic reasoning required to navigate the full 17k+ US-GAAP taxonomy and are poorly aligned with LLM capabilities. In parallel, financial document understanding has advanced from entity and relation extraction (Shah et al., 2023; Sharma et al., 2022) to joint text–table numerical reasoning in QA benchmarks such as FinQA, TAT-QA, and DocVQA (Chen et al., 2021; Zhu et al., 2021; Mathew et al., 2021), yet these tasks do not require alignment to standardized accounting concepts. Recent LLM-based tools such as XBRL-Agent (Han et al., 2024) explore filing analysis but stop short of end-to-end fact-to-taxonomy evaluation. In contrast, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: reframes XBRL tagging as a two-stage extract-and-align task, explicitly targeting multi-modal fact extraction and fine-grained alignment to official US-GAAP definitions.
3. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
3.1. Task Formulation
@@
Formally, given a financial document $D$ containing text and structured tables, a set of predefined XBRL value types $L$, and a taxonomy $\mathcal{T}$ of financial concepts, the XBRL tagging task is to identify all numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. Here, Fact denotes the extracted numerical value as expressed in context, Type specifies its XBRL data type in $L$, and Tag refers to the linked concept in $\mathcal{T}$ (e.g., us-gaap:CashAndDueFromBanks). Each numerical fact is assigned a single, most specific taxonomy concept. The value type set $L$ includes monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType, corresponding to monetary amounts, ratios, share counts, per-share values, and integer quantities, respectively. We formalize the task as a mapping:
(1) $f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$
@@
Inspired by information extraction and alignment works (Wu et al., 2019; Wadden et al., 2019; Mukherjee et al., 2022; Wang et al., 2022) and after discussing with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
3.1.1. Financial Numeric Identification (FinNI)
The first subtask of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: focuses on identifying numerical values in a financial document and assigning each a coarse-grained value data type. This corresponds to detecting the Fact and Type components of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task. We formalize this subtask as a mapping:
(2) $f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$